OpenStudy (turtwig):

solve for c a(b-c)=d I'm very confused. I would appreciate if someone could explain each step to me.

OpenStudy (mecharv):

Hey there :)

OpenStudy (mecharv):

So we have to get c on the LHS, \[a(b-c)=d\] \[ab-ac=d\] Bringing ac to the LHS, \[ac=ab-d\] Dividing the equation by a to have only c in the LHS \[c=\frac{ ab-d }{ a }\] \[c=b-\frac{ d }{ a }\]

OpenStudy (mecharv):

@turtwig Does that make sense?

OpenStudy (mecharv):

@turtwig I am planning to leave once you understand this... so.. are you stuck anywhere?

OpenStudy (turtwig):

Hi @mecharv , sorry I'm having some technical difficulties. I don't understand what you did to bringing ac to the LHS. If I have ab-ac=d, what makes it ac=ab-d? Where did the negative sign in ad-ac go?

OpenStudy (mecharv):

See carefully, I have taken ac to the RHS and d to the LHS And then I have switched LHS and RHS for convinience... So negative ac becomes positive when it goes to RHS and the positive d gets a negative sign... Is this clear?

OpenStudy (retireed):

ab−ac=d Bringing ac to the LHS, ac=ab−d It is a little confusing, I agree. He basically did two steps in one move.... I will show you both ab−ac=d -ac = d - ab ab subtracted from both ac = -d + ab multiply both sides by -1 ac = ab - d just reorganizing the RHS Is that clearer?

OpenStudy (turtwig):

Oh ok I think I see what you guys did. But what is the reason for multiplying each side by -1? I understand it makes -ac into a positive, but is that the only reason? @mecharv @retirEEd

OpenStudy (retireed):

yeah otherwise you would would to divided both sides by -a, which is doable, but a little more difficult to understand

OpenStudy (turtwig):

@retirEEd @mecharv Thank you both very much!

OpenStudy (retireed):

I did very little, but your welcomed.

OpenStudy (mecharv):

@retirEEd Woah you took a lot of pain indeed :O @turtwig Sorry for the confusion :( My bad.

OpenStudy (turtwig):

Don't apologize @mecharv , I'm very grateful for your clear answers to this and my other question! I think you have over-estimated my intelligence is all x)